OpenStudy (anonymous):

the measure of one interior angle of a REGULAR polygon is sixty degrees more than twice the measure of an exterior angle at the same vertex. how many sides has this polygon have?

OpenStudy (callisto):

let x be the measure of one interior angle. '' the measure of one interior angle of a REGULAR polygon is sixty degrees more than twice the measure of an exterior angle at the same vertex. '' => exterior angle = (x-60) /2 x + [ (x-60) /2 ] = 180 (adj. angles on st. line) 2x + x - 60 = 180(2) Solve x. Let n no the number of sides of the polygon. By angle sum polygon, (n-2) (180) = n(x) Sub x you've calculated into the equation above and solve n.